OpenStudy (anonymous):

FAN AND MEDAL!!! Determine which system below will produce infinitely many solutions. A) -4x + 6y = 20 4x - 3y = -2 B) -2x + 4y = 18 4x - 8y = -36 C) 3x - 2y = 5 -6x - 4y = -10 D) 5x - 3y = 12 15x - 9y = 24

OpenStudy (triciaal):

to produce infinitely many solutions the 2 functions are actually the same one is a factor of the other

OpenStudy (anonymous):

so its B?

OpenStudy (triciaal):

yes!

OpenStudy (anonymous):

Thanks! can u help with another one please?

OpenStudy (triciaal):

post

OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -1 x = 0 x = 2 There is no x value, as there is no solution. x can be any value, as there is an infinite number of solutions.

OpenStudy (triciaal):

by substitution write an expression for one of the variable and substitute this expression in the second equation. you now have one equation with one unknown find the value. use this value in the expression to find the value of the other variable. good practice to check your answer by putting your solution in either of the originals.

OpenStudy (triciaal):

from eqn 1 y = 1-2x in eqn 2 4x + 2(1-2x) = -1 4x +2 -4x = -1

OpenStudy (triciaal):

no so no solution

OpenStudy (anonymous):

Thanks! I understand everything now!

OpenStudy (anonymous):

Can i tag u if i need more help please?

OpenStudy (triciaal):

great!! you are welcome